# Trivial Isomorphisms between Reduced Products

### v2

@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
Because of this and $\operatorname{{\mathsf{OCA}_{T}}}$, (1) of Proposition 3.6 holds. Fix ${\mathcal{F}}_{n}$, for $n\in{\mathbb{N}}$, such that ${\mathcal{F}}=\bigcup_{n}{\mathcal{F}}_{n}$, and for all $n$ and all $g,h$ in ${\mathcal{F}}_{n}$ we have $|\operatorname{Diff}(g,h)|\leq n$. Since ${\mathcal{J}}_{\textrm{cont}}$ is nonmeager, there is $n$ such that ${\mathcal{X}}_{n}=\{A\mid(A,g_{A})\in{\mathcal{F}}_{n}\}$ is nonmeager. Let
${\mathcal{G}}_{n}=\{(B,h_{B})\mid(\exists A\in{\mathcal{X}}_{n})\ B\setminus A\in\operatorname{Fin},\\ h_{B}(n)=g_{A}(n)\text{ for }n\in B\cap A,g_{A}(n)=\emptyset\text{ for }n\in B\setminus A\}.$
Then $C\mapsto h_{B}[C]$ is a lifting of $\Phi$ on ${\mathcal{P}}(B)$ for every $(B,h_{B})\in{\mathcal{G}}_{n}$. Let ${\mathcal{Z}}$ be the closure of ${\mathcal{G}}_{n}$ in ${\mathcal{P}}({\mathbb{N}})\times\operatorname{Fin}^{\mathbb{N}}$. The set $\{B\mid(\exists A\in{\mathcal{X}}_{n})B\setminus A\in\operatorname{Fin}\}$ is closed under finite modifications of its elements and it is therefore dense in the standard topology of ${\mathcal{P}}({\mathbb{N}})$. Therefore for every $B\in{\mathcal{P}}({\mathbb{N}})$ the section ${\mathcal{Z}}_{B}=\{h\mid(B,h)\in{\mathcal{G}}_{n}\}$ is nonempty. By the Jankov–von Neumann uniformization theorem (Theorem A.1) there is a C-measurable function $\Theta\colon{\mathcal{P}}({\mathbb{N}})\to\operatorname{Fin}^{\mathbb{N}}$ such that $(A,\Theta(A))\in{\mathcal{Z}}$ for all $A$.
@@
We claim that $\operatorname{Diff}(\Theta(A),g_{A})$ is finite for every $A\in{\mathcal{J}}_{\textrm{cont}}$. Assume otherwise. Since ${\mathcal{X}}_{n}$ is nonmeager, its hereditary closure $\hat{\mathcal{X}}_{n}=\{A\subseteq{\mathbb{N}}\mid(\exists C\in{\mathcal{F}}_{n})A\subseteq C\}$ is nonmeager and hereditary and by [8, Theorem 3.10.1] there is an infinite $B\in\hat{\mathcal{X}}_{n}$ included in $\operatorname{Diff}(\Theta(A),g_{A})$. Then $\operatorname{Diff}(\Theta(A),g_{A})\subseteq\operatorname{Diff}(g_{A},h_{B})\cup\operatorname{Diff}(h_{B},\Theta(A))$, which is the union of two finite sets; contradiction.
@@
$\tilde{\Theta}(A)=\bigcup\Theta_{A}[A]$
is a C-measurable lifting of $\Phi$ on ${\mathcal{J}}_{\textrm{cont}}$. By [8, Lemma 3.3.4], $\Phi$ decomposes as $\Phi_{1}\oplus\Phi_{2}$ where $\Phi_{1}$ has a continuous lifting and the kernel of $\Phi_{2}$ is nonmeager. By [38], $\Phi_{1}$ has a completely additive lifting, as required. ∎
@@ 5. Proof of Theorem 7
There are only countably many choices for $\mathbb{I}(p)$. We prove that if $\mathbb{I}(p)=\mathbb{I}(q)$, then $p$ and $q$ are compatible. This concludes the proof of the claim.
Define $F_{r}=F_{p}\cup F_{q}$ and choose $n_{r}$ large enough such that $A\cap B\subseteq n_{r}$ for every $(a,A),(b,B)\in F_{r}$ with $(a,A)\neq(b,B)$. It remains to define $s^{0}_{r}(n)$ and $s^{1}_{r}(n)$ for every $n\in n_{r}\setminus n_{p}$ in such a way that $r\in\mathbb{P}$ with $r\leq p,q$. If such $n$ has the property that there exists at most one $(a,A)\in F_{r}$ such that $n\in A$, then $s^{0}_{r}(n)$ and $s^{1}_{r}(n)$ can be defined arbitrarily subject only to $a(n)\in\{s^{0}_{p}(n),s^{1}_{p}(n)\}$. Else, if $n\in n_{r}\setminus n_{p}$ is such that the set $\{(c,C)\in F_{r}:n\in C\}$ has at least two elements, then since $p$ and $q$ are both $\mathbb{P}$-conditions, there exist $(a,A)\in F_{p}$ and $(b,B)\in F_{q}$ such that
$\{(c,C)\in F_{r}:n\in C\}=\{(a,A),(b,B)\},$
in this case define $s^{0}_{r}(n)=a(n)$ and $s^{1}_{r}(n)=b(n)$. ∎
Let $\{(a_{\xi},A_{\xi}):\xi<\omega_{1}\}$ be a one-to-one enumeration of $H$ (by having passed to a subset if necessary, we may assume that $H$ has cardinality $\aleph_{1}$). Consider the sets $E_{\xi}=\{q\mid(\exists\eta>\xi)(a_{\eta},A_{\eta})\in F_{q}\}$, for $\xi<\omega_{1}$, and $E_{n}=\{q\mid n_{q}>n\}$, for $n\in{\mathbb{N}}$. By the earlier mentioned standard property of ccc posets, there exists a condition $p$ such that these sets are all dense below $p$, so by $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}(\sigma$-linked) there exists a filter $G\subseteq\mathbb{P}$ that meets them all. Then $\bigcup_{q\in G}F_{q}\subseteq H$ is uncountable and the set $\{n_{q}:q\in G\}$ is infinite. Hence, by shrinking $H$ if necessary we can actually assume that there exist $s^{0},s^{1}\in\prod_{n}{\mathcal{M}}_{n}$ such that
@@
This proves that ${\mathcal{J}}_{\rm{prod}}$ intersects ${\mathcal{A}}$ nontrivially. By applying this to an uncountable subfamily of ${\mathcal{A}}$, we conclude that ${\mathcal{J}}_{\rm{prod}}$ includes all but countably many elements of ${\mathcal{A}}$. ∎
@@ 5.2. Uniformization
Consider the Polish space ${\mathcal{Y}}_{n}=({\mathcal{N}}_{n})^{{\mathcal{M}}_{n}}$. Then $\{g^{A}\mid A\in{\mathcal{J}}_{\rm{prod}}\}$ is a family of partial selectors as defined and studied in §3.1. The following shows that the functions $g^{A}$, for $A\in{\mathcal{J}}_{\rm{prod}}$, moreover form a coherent family in the sense of [8, §2].
@@
Since ${\mathcal{J}}_{\rm{prod}}$ is nonmeager, Proposition 5.2 implies that there is $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\rm{prod}}$. For $x\in Z$ let $n(x)=\max(\operatorname{Diff}(g^{f(x)},g^{B}))$ (with $\max\emptyset=0$). Since $Z$ is uncountable, there is $n$ such that $Z^{\prime}=\{x\in Z\mid n(x)<n\}$ is uncountable. Choose distinct $y,z$ in $Z^{\prime}$ satisfying $\Delta(y,z)\geq n$ and fix $j\in(\operatorname{Diff}(g^{f(y)},g^{f(z)})\cap B)\setminus n$. Then $g^{B}_{j}=g^{f(y)}_{j}\neq g^{f(z)}_{j}=g^{B}_{j}$; contradiction. ∎
@@
${\mathcal{Y}}_{a}=\{b\in\prod_{n}{\mathcal{N}}_{n}\mid(a,b)\in{\mathcal{Y}}^{\prime}\}$
@@ 5.3. Automorphisms with Baire-measurable liftings
Consider first the case when all ${\mathcal{M}}_{n}$ are finite. In this case we will not need $\operatorname{{\mathsf{OCA}_{T}}}$. Since $\Theta$ is Baire-measurable, it is continuous on a dense $G_{\delta}$ set $G$. By the standard representation of dense $G_{\delta}$ sets in $\prod_{n}{\mathcal{M}}_{n}$ (see e.g., [12, Theorem 9.9.1]) there are a partition ${\mathbb{N}}=S_{0}\sqcup S_{1}$, $a_{0}\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$, and $a_{1}\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$ such that $\{b\in\prod_{n}{\mathcal{M}}_{n}\mid(\forall^{\infty}n\in S_{j})\,b(n)=a_{j}(n)\}\subseteq G$ for $j=0,1$. (This is precisely where the finiteness of ${\mathcal{M}}_{n}$ gets used, as this characterisation does not hold for infinite sets). If $x\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$ and $y\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$, let $x^{\frown}y$ be the concatenation of $x$ and $z$, that is, the unique $z\in\prod_{n}{\mathcal{M}}_{n}$ which equals $x$ on $S_{0}$ and $y$ on $S_{1}$.
@@
ordered by the end-extension. By the assumption on $\Upsilon$, in $T$ there is no branch $t_{i}=(u_{i},v_{i},s_{i})$ such that $\bigcup s_{i}$ is infinite. Fix a node $t=(u,v,s)\in T$ with the property that if $t^{\prime}=(u^{\prime},v^{\prime},s^{\prime})\in T$ extends $t$ then $s^{\prime}\setminus s$ is empty, and let $m$ be such that $u\in\prod_{j<m}{\mathcal{M}}_{j}$.
We claim that for all $a$ and $b$ in $[u]$ and all $j\geq m$, if $a(j)=b(j)$ then $\Upsilon(a)(j)=\Upsilon(b)(j)$. Assume otherwise and fix $c\in[v]$ such that $c(j)=a(j)$. Then $\Upsilon(a)(j)=\Upsilon(c)(j)=\Upsilon(b)(j)$; contradiction.
@@
Suppose that ${\mathcal{X}}$ has an uncountable $K_{0}$-homogeneous set, ${\mathcal{H}}$. By Claim 5.11 there is $G\in{\mathcal{X}}$ such that =${\mathcal{H}}’=\{F\in{\mathcal{H}}\mid F_{n}\subseteq G_{n}$ for all $n\}$ is uncountable. For every $F\in{\mathcal{X}}$ there is $m(F)$ such that for all $j\geq m(F)$ we have $h^{G}_{j}\restriction F_{j}=h^{F}_{j}$. Fix $m$ such that ′′=${\mathcal{H}}^{\prime\prime}=\{F\in{\mathcal{H}}^{\prime}\mid m(F)=m\}$ is uncountable. Since the set $\bar{G}=\prod_{j<m}G_{j}$ is finite, there are only countably many possibilities for the restriction of $h$ to this set. Therefore there are $F$ and $F^{\prime}$ in ${\mathcal{H}}^{\prime\prime}$ such that $h_{j}^{F}=h_{j}^{F^{\prime}}$ for all $j<m$. This implies $\{F,F^{\prime}\}\in K_{1}$; contradiction.

### v3

@@ 4. Automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$
Because of this and $\operatorname{{\mathsf{OCA}_{T}}}$, (1) of Proposition 3.7 holds. Fix ${\mathcal{F}}_{n}$, for $n\in{\mathbb{N}}$, such that ${\mathcal{F}}=\bigcup_{n}{\mathcal{F}}_{n}$, and for all $n$ and all $(g_{A},A)$ and $(g_{B},B)$ in ${\mathcal{F}}_{n}$ we have $|\operatorname{Diff}(g_{A},g_{B})|\leq n$. Since ${\mathcal{J}}_{\textrm{cont}}$ is nonmeager, there is $n$ such that
${\mathcal{X}}_{n}=\{A\mid(g_{A},A)\in{\mathcal{F}}_{n}\}$
is nonmeager. Let
${\mathcal{G}}=\overline{\{(g,B)\mid\exists(g_{A},A)\in{\mathcal{F}}_{n}(B\subseteq A,g=g_{A}\restriction B)\}}.$
The closure is taken in the space of partial selectors as in Definition 3.4. Since each ${\mathcal{Y}}_{n}$ is the discrete space $\operatorname{Fin}$, this is a Polish topology. Therefore the set
$\mathcal{W}=\{B\mid(\exists g)(g,B)\in{\mathcal{G}}\}$
is, being the projection of a closed set, an analytic subset of ${\mathcal{P}}({\mathbb{N}})$. It is also hereditary and (because it includes ${\mathcal{X}}_{n}$) nonmeager. Note also that ${\mathcal{G}}$ has the critical property that for all $(g_{0},B_{0})$ and $(g_{1},B_{1})$ in ${\mathcal{G}}$ we have $|\operatorname{Diff}(g_{0},g_{1})|\leq n$.
Since $\mathcal{W}$ is nonmeager and analytic, there is a basic open subset $U$ of ${\mathcal{P}}({\mathbb{N}})$ such that $U\setminus\mathcal{W}$ is meager (see e.g. [26, Proposition 8.26]). It is of the form $U=\{A\mid A\cap J_{0}=s_{0}\}$ for some finite interval $J_{0}$ and $s_{0}\subseteq J_{0}$. As $\mathcal{W}$ is hereditary, we may assume that $s_{0}=\emptyset$.
Since $\mathcal{W}$ is relatively comeager in ${\mathcal{P}}({\mathbb{N}}\setminus J_{0})$, there is a partition of ${\mathbb{N}}\setminus J_{0}$ into finite intervals $J_{i}$, for $i\geq 1$, and $s_{i}\subseteq J_{i}$ for all $i$ such that
$\{A\subseteq{\mathbb{N}}\setminus J_{0}\mid(\exists^{\infty}i)A\cap J_{i}=s_{i}\}\subseteq\mathcal{W}.$
Let $S=\bigcup s_{i}$. As $s_{i}\subseteq J_{i}$ and $\{J_{i}\}$ partitions $\mathbb{N}\setminus J_{0}$, $S\in\mathcal{W}$.
Let $A_{\textrm{even}}=\bigcup_{i\geq 1,i\ \textrm{even}}J_{i}$, $A_{\textrm{odd}}=\bigcup_{i\ \text{odd}}J_{i}$, $S_{\textrm{even}}=S\cap A_{\textrm{even}}$, and $S_{\textrm{odd}}=S\cap A_{\textrm{odd}}$, we have that for every $B\subseteq{\mathbb{N}}$, both $B_{\textrm{even}}=(B\cap A_{\textrm{even}})\cup S_{\textrm{odd}}$ and $B_{\textrm{odd}}=(B\cap A_{\textrm{odd}})\cup S_{\textrm{even}}$ belong to $\mathcal{W}$. Using this notation, let
$\begin{split}{\mathcal{V}}=\{&(g,B)\mid B\subseteq{\mathbb{N}},g:B\to\operatorname{Fin},\\ &\text{and some $h_{\textrm{even}}$ and $h_{\textrm{odd}}$ satisfy }(h_{\textrm{even}},B_{\textrm{even}})\in{\mathcal{G}},(h_{\textrm{odd}},B_{\textrm{odd}})\in{\mathcal{G}},\\ &B\subseteq B_{\textrm{even}}\cup B_{\textrm{odd}},\operatorname{Diff}(h_{\textrm{even}},g)=\emptyset\text{ and }|\operatorname{Diff}(h_{\textrm{odd}},g)|\leq n\}.\end{split}$
Claim 4.8.
For every $B\subseteq{\mathbb{N}}$ there is $g$ such that $(g,B)\in{\mathcal{V}}$ and every such $g$ satisfies $|\operatorname{Diff}(g,g_{C})|\leq 3n$ for every $C\in{\mathcal{X}}_{n}$.
Proof.
Fix $B\subseteq{\mathbb{N}}$. Since both $B_{\textrm{even}}$ and $B_{\textrm{odd}}$ belong to $\mathcal{W}$, some $h_{\textrm{even}}$ and $h_{\textrm{odd}}$ satisfy $(h_{\textrm{even}},B_{\textrm{even}})\in{\mathcal{G}}$ and $(h_{\textrm{odd}},B_{\textrm{odd}})\in{\mathcal{G}}$. Let
$g=h_{\textrm{even}}\restriction(B_{\textrm{even}}\cap B)\cup(h_{\textrm{odd}}\restriction(B\setminus B_{\textrm{even}})).$
By the homogeneity of the set ${\mathcal{G}}$ we have $|\operatorname{Diff}(h_{\textrm{odd}},h_{\textrm{even}})|\leq n$, and therefore $\operatorname{Diff}(h_{\textrm{odd}},g)\subseteq\operatorname{Diff}(h_{\textrm{odd}},h_{\textrm{even}})$ implies $(g,B)\in{\mathcal{V}}$.
To prove the second part of the claim, fix $C\in\mathcal{X}_{n}$. Then $(g_{C},C)\in\mathcal{G}$, and therefore $\max(|\operatorname{Diff}(g_{C},h_{\textrm{even}})|,|\operatorname{Diff}(g_{C},h_{\textrm{odd}})|)\leq n$. Since
$\operatorname{Diff}(g,g_{C})\subseteq$ $\operatorname{Diff}(g\restriction B_{\textrm{even}},h_{\textrm{even}})\cup\operatorname{Diff}(g\restriction B_{\textrm{odd}},h_{\textrm{odd}})$
$\cup\operatorname{Diff}(h_{\textrm{even}},g_{C}\restriction B_{\textrm{even}})\cup\operatorname{Diff}(h_{\textrm{odd}},g_{C}\restriction B_{\textrm{odd}}),$
and $\operatorname{Diff}(g\restriction B_{\textrm{even}},h_{\textrm{even}})=\emptyset$, we have $|\operatorname{Diff}(g,g_{C})|\leq 3n$ as required. ∎
By the Jankov–von Neumann uniformization theorem (Theorem A.1) there is a C-measurable function $\Theta\colon{\mathcal{P}}({\mathbb{N}})\to\bigcup_{A\subseteq{\mathbb{N}}}\prod_{n\in A}{\mathcal{Y}}_{n}$ such that $(\Theta(A),A)\in{\mathcal{V}}$ for all $A$.
@@
We claim that $\operatorname{Diff}(\Theta(A),g_{A})$ is finite for every $A\in{\mathcal{J}}_{\textrm{cont}}$. Assume otherwise and fix $A\in{\mathcal{J}}_{\textrm{cont}}$ for which this set is infinite. Since ${\mathcal{X}}_{n}$ is nonmeager, its hereditary closure $\hat{\mathcal{X}}_{n}=\{B\subseteq{\mathbb{N}}\mid(\exists C\in{\mathcal{X}}_{n})B\subseteq C\}$ is nonmeager and hereditary and by [8, Theorem 3.10.1] there is an infinite $B\in\hat{\mathcal{X}}_{n}$ included in $\operatorname{Diff}(\Theta(A),g_{A})$. If $C\in{\mathcal{X}}_{n}$ with $B\subseteq C$, then $B\subseteq\operatorname{Diff}(g_{A},g_{C})\cup\operatorname{Diff}(g_{C},\Theta(A))$. However, each one of these two sets is finite: $\operatorname{Diff}(g_{A},g_{C})$ is finite because both $A$ and $C$ belong to ${\mathcal{J}}_{\textrm{cont}}$ and $|\operatorname{Diff}(g_{C},\Theta(A))|\leq 3n$ by Claim 4.8; contradiction.
@@
$\tilde{\Theta}(A)=\bigcup\Theta[A]$
is a C-measurable lifting of $\Phi$ on ${\mathcal{J}}_{\textrm{cont}}$. By [8, Lemma 3.3.4], $\Phi$ decomposes as $\Phi_{1}\oplus\Phi_{2}$ where $\Phi_{1}$ has a continuous lifting and the kernel of $\Phi_{2}$ is nonmeager. By [8, Theorem 1.6.1], $\Phi_{1}$ has a completely additive lifting, as required. ∎
@@ 5. Proof of Theorem 7
There are only countably many choices for $\mathbb{I}(p)$. We prove that if $\mathbb{I}(p)=\mathbb{I}(q)$, then $p$ and $q$ are compatible. This will conclude the proof of the claim.
Define $F_{r}=F_{p}\cup F_{q}$ and choose $n_{r}$ large enough to have $A\cap B\subseteq n_{r}$ for all distinct $(a,A)$ and $(b,B)$ in $F_{r}$. It remains to define $s^{0}_{r}(n)$ and $s^{1}_{r}(n)$ for every $n\in n_{r}\setminus n_{p}$ so that $r$ is a condition. Any such $r$ will extend both $p$ and $q$.
Fix $n\in n_{r}\setminus n_{p}$. Since $p\in\mathbb{P}$, there is at most one $(a,A)\in F_{p}$ such that $n\in A$. Similarly, there is at most one $(b,B)\in F_{q}$ such that $n\in B$. If both $(a,A)$ and $(b,B)$ are defined, let $s^{0}_{r}(n)=a(n)$ and $s^{1}_{r}(n)=b(n)$. Otherwise, we may choose one (or both) $s^{j}_{r}(n)$ arbitrarily.
Clearly $r$ defined in this manner belongs to $\mathbb{P}$ and extends both $p$ and $q$. ∎
Let $\{(a_{\xi},A_{\xi}):\xi<\omega_{1}\}$ be a one-to-one enumeration of $H$ (by having passed to a subset if necessary, we may assume that $H$ has cardinality $\aleph_{1}$). Consider the sets $E_{\xi}=\{q\in\mathbb{P}\mid(\exists\eta>\xi)(a_{\eta},A_{\eta})\in F_{q}\}$, for $\xi<\omega_{1}$, and $E_{n}=\{q\in\mathbb{P}\mid n_{q}>n\}$, for $n\in{\mathbb{N}}$. Each $E_{n}$ is dense in $\mathbb{P}$, as can be seen from the proof of Claim 5.3. Each one of the $E_{\xi}$ is a nonempty subset of $\mathbb{P}$. Since $\mathbb{P}$ has the countable chain condition, some $p\in\mathbb{P}$ forces that the generic filter $\dot{G}$ intersects uncountably many of the $E_{\xi}$ nontrivially. (To see this, assume otherwise and fix a maximal antichain consisting of conditions that decide the supremum of the set of all $\xi$ such that $\dot{G}\cap E_{\xi}\neq\emptyset$. Since the antichain is countable we have $E_{\xi}=\emptyset$ for all but countably many $\xi$; contradiction.) Since $\xi<\eta$ implies $E_{\xi}\supseteq E_{\eta}$, each $E_{\xi}$ is dense below $p$.
By $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}(\sigma$-linked) there exists a filter $G\subseteq\mathbb{P}$ that intersects all $E_{n}$ and all $E_{\xi}$ nontrivially. Therefore $\bigcup_{q\in G}F_{q}\subseteq H$ is uncountable and the set $\{n_{q}:q\in G\}$ is infinite. Hence, by shrinking $H$ if necessary we can actually assume that there exist $s^{0},s^{1}\in\prod_{n}{\mathcal{M}}_{n}$ such that
@@
This proves that ${\mathcal{J}}_{\rm{prod}}$ intersects ${\mathcal{A}}$ nontrivially. Since every subfamily of $\mathcal{A}$ is tree-like, and therefore if it is in addition uncountable then ${\mathcal{J}}_{\rm{prod}}$ intersects it, ${\mathcal{J}}_{\rm{prod}}$ includes all but countably many elements of ${\mathcal{A}}$. ∎
@@ 5.2. Uniformization
Consider the Polish space ${\mathcal{Y}}_{n}=({\mathcal{N}}_{n})^{{\mathcal{M}}_{n}}$. Then $\{(g^{A},A)\mid A\in{\mathcal{J}}_{\rm{prod}}\}$ is a family of partial selectors as defined and studied in §3.1. The following shows that the functions $g^{A}$, for $A\in{\mathcal{J}}_{\rm{prod}}$, moreover form a coherent family in the sense of [8, §2].
@@
Proposition 5.2 implies that ${\mathcal{J}}_{\rm{prod}}$ is nonmeager (recall that this conclusion is included in the assumptions of Proposition 5.6), and (again by Talagrand and Jalali-Naini, [8, Theorem 3.10.1]) it intersects every perfect tree-like almost disjoint family. Fix $B\in{\mathcal{A}}\cap{\mathcal{J}}_{\rm{prod}}$. For $x\in Z$ let $n(x)=\max(\operatorname{Diff}(g^{f(x)},g^{B}))$ (with $\max\emptyset=0$). Since $Z$ is uncountable, there is $n$ such that $Z^{\prime}=\{x\in Z\mid n(x)<n\}$ is uncountable. Choose distinct $y,z$ in $Z^{\prime}$ satisfying $\Delta(y,z)\geq n$ and fix $j$ in $(\operatorname{Diff}(g^{f(y)},g^{f(z)})\cap B)\setminus n$. Then $g^{B}_{j}=g^{f(y)}_{j}\neq g^{f(z)}_{j}=g^{B}_{j}$; contradiction. ∎
@@
${\mathcal{Y}}_{a}=\{b\in\prod_{n}{\mathcal{N}}_{n}\mid(a,b)\in{\mathcal{Y}}\}$
@@ 5.3. Automorphisms with Baire-measurable liftings
Consider first the case when all ${\mathcal{M}}_{n}$ are finite. In this case we will not need $\operatorname{{\mathsf{OCA}_{T}}}$. Since $\Theta$ is Baire-measurable, it is continuous on a dense $G_{\delta}$ set $G$. By the standard representation of dense $G_{\delta}$ sets in $\prod_{n}{\mathcal{M}}_{n}$ (see e.g., [12, Theorem 9.9.1]) there are a partition ${\mathbb{N}}=S_{0}\sqcup S_{1}$, $a_{0}\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$, and $a_{1}\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$ such that for $j\in\{0,1\}$ we have $\{b\in\prod_{n}{\mathcal{M}}_{n}\mid b\restriction S_{j}=a_{j}\}\subseteq G$. (This is precisely where the finiteness of ${\mathcal{M}}_{n}$ gets used, as this characterisation does not hold for infinite sets). If $x\in\prod_{n\in S_{0}}{\mathcal{M}}_{n}$ and $y\in\prod_{n\in S_{1}}{\mathcal{M}}_{n}$, let $x^{\frown}y$ be the concatenation of $x$ and $y$, that is, the unique $z\in\prod_{n}{\mathcal{M}}_{n}$ which equals $x$ on $S_{0}$ and $y$ on $S_{1}$.
@@
ordered by the end-extension. By the assumption on $\Upsilon$, in $T$ there is no branch $t_{i}=(u_{i},v_{i},s_{i})$ such that $\bigcup s_{i}$ is infinite. Fix a node $t=(u,v,s)\in T$ with the property that if $t^{\prime}=(u^{\prime},v^{\prime},s^{\prime})\in T$ extends $t$ then $s^{\prime}\setminus s$ is empty, and let $m$ be such that $u\in\prod_{j<m}{\mathcal{M}}_{j}$. We claim that for all $a$ and $b$ in $[u]$ and all $j\geq m$, if $a(j)=b(j)$ then $\Upsilon(a)(j)=\Upsilon(b)(j)$. Assume otherwise and fix $c\in[v]$ such that $c(j)=a(j)$. Then $\Upsilon(a)(j)=\Upsilon(c)(j)=\Upsilon(b)(j)$; contradiction.
@@
Suppose that ${\mathcal{X}}$ has an uncountable $K_{0}$-homogeneous set, ${\mathcal{H}}$. By Claim 5.11 there is $G\in{\mathcal{X}}$ such that =${\mathcal{H}}’=\{F\in{\mathcal{H}}\mid F_{n}\subseteq G_{n}$ for all $n\}$ is uncountable. For every $F\in{\mathcal{H}}$ there is $m(F)$ such that for all $j\geq m(F)$ we have $h^{G}_{j}\restriction F_{j}=h^{F}_{j}$. Fix $m$ such that ′′=${\mathcal{H}}^{\prime\prime}=\{F\in{\mathcal{H}}^{\prime}\mid m(F)=m\}$ is uncountable. Since the set $\bar{G}=\prod_{j<m}G_{j}$ is finite, there are only countably many possibilities for the restriction of $h$ to this set. Therefore there are $F$ and $F^{\prime}$ in ${\mathcal{H}}^{\prime\prime}$ such that $h_{j}^{F}=h_{j}^{F^{\prime}}$ for all $j<m$. This implies $\{F,F^{\prime}\}\in K_{1}$; contradiction.